A sample of gas in a balloon has an initial temperature of 22 ∘C and a volume of 1180 L . If the temperature changes to 57 ∘C , and there is no change of pressure or amount of gas, what is the new volume, V2, of the gas?

What Celsius temperature, T2, is required to change the volume of the gas sample in Part A (T1 = 22 ∘C , V1= 1180 L ) to a volume of 2360 L ? Assume no change in pressure or the amount of gas in the balloon.
Answer #1

Initial        Final

T1 = 22+273 = 295K    T2 = 57+273 = 330K

V1 = 1180L V2 =

V1/T1    = V2/T2

V2    = V1T2/T1

   = 1180*330/295                         = 1320L

The new V2 volume of gas = 1320L

2.

Initial        Final

T1 = 22+273 = 295K    T2 =

V1 = 1180L V2 = 2360L

V1/T1    = V2/T2

T2   = V2T1/V1

                           = 2360*295/1180   = 590K

                             = 590-273   = 3170C

Celsius temperature, T2   = 3170C >>>>answer